OpenStudy (anonymous):

solve 5/2x+1/4x=5/4+x what is the solution x=

OpenStudy (anonymous):

is that (5/4)+(x) or 5/(4+x)?

jhonyy9 (jhonyy9):

10(4+x) +(4+x)=20x 40 +10x +4 +x=20x 44=9x x=44/9

OpenStudy (anonymous):

is that simplified though

jhonyy9 (jhonyy9):

how ?

OpenStudy (anonymous):

wouldnt it be 5/7

jhonyy9 (jhonyy9):

how you have got it ?